# MLLM as Retriever: Interactively Learning Multimodal Retrieval for Embodied Agents

### v1

@@ Abstract
MLLM agents demonstrate potential for complex embodied tasks by retrieving multimodal task-relevant trajectory data. However, current retrieval methods primarily focus on surface-level similarities of textual or visual cues in trajectories, neglecting their effectiveness for the specific task at hand. To address this issue, we propose a novel method, MLLM As ReTriever (MART), which enhances the performance of embodied agents by utilizing interaction data to fine-tune an MLLM retriever based on preference learning, such that the retriever fully considers the effectiveness of trajectories and prioritize them for unseen tasks. We also introduce Trajectory Abstraction, a mechanism that leverages MLLMs’ summarization capabilities to represent trajectories with fewer tokens while preserving key information, enabling agents to better comprehend milestones in the trajectory. Experimental results across various environments demonstrate our method significantly improves task success rates in unseen scenes compared to baseline methods. This work presents a new paradigm for multimodal retrieval in embodied agents, by fine-tuning a general-purpose MLLM as the retriever to assess trajectory effectiveness. All benchmark task sets and simulator code modifications for action and observation spaces will be released.
@@ 4.1 Experimental Setup
Table 1: Environment complexity comparison. AI2-THOR contains more than four times the number of interactive objects per scene, and more complex object hierarchy, compared to LEGENT. E.g., AI2-THOR supports relationships such as “inside" (e.g., inside a microwave (Figure 7(a)), or inside an open container, like a sink (Figure 7(c))).
@@ 4.1.1 Environments
To validate the effectiveness of our method in various environments, we perform evaluations on multiple scenarios in two environments, AI2-THOR (Kolve et al., 2017) and LEGENT (Cheng et al., 2024). Unlike LLM agents, which use only text as input and simplify the action space by employing teleportation actions (e.g., “move directly to target"), we believe that MLLM agents should undertake more challenging tasks as they have access to visual input and are no longer limited to a ‘blind’ mode of operation. Therefore, both target environments are multimodal, whose observations are egocentric images, and both make use of fine-grained control actions.
@@
Table 4 and Figure 4 show a comparison between the two environments. Besides having many more objects in its scenes, AI2-THOR supports more complex object hierarchy relationships, such as “inside" (e.g., “inside a cabinet", which require open/close interaction to complete a task, or “a cup inside a sink", i.e., an open container), not supported in LEGENT  [^1] Although there are cabinets in LEGENT, they do not need to be opened/closed to complete tasks..
@@ 4.1.2 Task Settings
AI2-THOR. To better assess the fine-grained control ability of MLLM agents to complete real-world embodied tasks, we integrate characteristics of two AI2-THOR-based benchmarks – ALFWorld (Shridhar et al., 2021) and ALFRed (Shridhar et al., 2020) – and built an environment setting that is more suitable for the MLLM agent. Since tasks are long-horizon, we follow the method in ALFRed and apply task decomposition to divide them into sub-tasks before execution. Each sub-task is then provided to the agent, and it determines sub-task success based on environmental feedback by itself (details in Appendix B.2 and  B.3 ). Once a sub-task is successfully completed, the agent proceeds to the next sub-task. Task types include pick_and_place, pick_clean_then_place, pick_cool_then_place, and pick_heat_then_place. Completing these tasks requires dozens of steps of navigation, as well as interaction with objects. There are 45 tasks comprising a total of 260 sub-tasks in training set, and 28 tasks including 158 sub-tasks in testing set.
@@ 4.1.4 Task Evaluation
Average Steps represents the average number of steps the agent takes to complete a task. In AI2-THOR  it indicates the number of steps to complete a full task, and AS-Sub represents the step average to complete a sub-task.
@@ 4.1.5 Baselines
LLaVA-Plain (LP) is a pre-trained LLaVA with no modified head and no finetuning. We use the probability of special token generation to represent the score. Its input is the same as MART, and it is prompted to output only Yes/No tokens, and the final score is calculated based on the probability of token generation (more details and limitations in Appendix A.3).
@@ 4.5 Case Study
We present two case studies for more in-depth discussion of MART’s capabilities handling challenges of the MLLM agent setting. The first case (Figure 5) demonstrates the effective handling of a very long-horizon trajectory. Given a 73-step task trajectory – “navigate to DishSponge, put it in the Plate, and place the Plate on the Cabinet" – Trajectory Abstraction identifies 5 key milestones. MART achieves an 80% success rate with an average of 28 steps, while both the similarity-based method and the agent without memory reach only 40% success rate, averaging 69.6 and 74.6 steps, respectively.
@@ 5 Conclusion and Limitations
As the context window is limited for the MLLM we used, it cannot accurately process a large number of images, so we only retrieve one trajectory at a time to input to the agent in the manner of one-shot learning. If multiple trajectories can be used as input at a time, as in few-shot learning, an agent could splice the skills corresponding to different sub-trajectories, which has the potential to further improve performance in more complex and longer-horizon tasks. Such setting, and also methods for retrieving trajectory sub-segments instead of full trajectories, are left as future work.
Appendix A Experimental setup details
@@
In this appendix we provide more low-level details on the implementation of MART experiments.
A.1 Multiple image input in LLaVA
@@
We utilized LLaVA’s single-image perception capabilities to enhance its multi-image perception. Specifically, we sampled images collected from the environments and employed the pre-trained LLaVA to describe the content of each image. We organized the responses into multi-image captioning data, which we then used to fine-tune LLaVA. After our verification, the fine-tuned LLaVA demonstrated the ability to perceive multiple input images. After fine-tuning with multiple image datasets, LLaVA serves as the basez model for our subsequent retriever training. We will release all the code.
A.2 Parameter settings
@@
A.3 Details of LLava-Plain
Building on existing work (Asai et al., 2024) (Sun et al., 2023b) that employs LLM for text retrieval, we use the generation probability of a special token to represent the score for LLaVA-Plain. In detail, the effectiveness score $s_{i}$ is measured by the probability of LLaVA-Plain to generate the special token ‘Yes’ and ‘No’, as in Equation 2, where $p\left(\mathrm{Yes/No}\right)$ denoted the probability of LLaVA-Plain to generate Yes or No. E.
@@ A.3 Details of LLava-Plain
In detail, the prompts we use for LLaVA-Plain and MART are presented in prompt 1 of Appendix E.
Appendix B AI2-THOR environment specifics
B.1 Setting differences
@@
ALFWorld adapts ALFRED for LLM agents (i.e., text-only) by simplifying it. Firstly, it provides text feedback as observation, detailing objects in the agent’s field of view along with their corresponding IDs. Secondly, it simplifies the action space by replacing all navigation actions with the teleportation action ‘go to’ and composite high-level actions like ‘heat’, ‘clean’, and ‘cool’, each involving multiple atomic interactions. For example, the “cool object a" action is equivalent to: ‘open the refrigerator’, ‘put object A inside the refrigerator’, ‘close the refrigerator’, ‘open the refrigerator’, and ‘pick up the object A’. These modifications significantly reduce task difficulty, while allowing LLM agents to perform in ALFWorld.
@@
All selected task sets and modified simulator code for adapting the MLLM agent will be released upon paper publication.
B.2 Task Decomposition
@@
B.3 Success Detection
In AI2-THOR, we provide the agent with the environment’s metadata and feedback in the Success Detection module. After each step is executed, the agent then determines whether the current sub-task is completed using a few-shot approach. For specific prompts, please refer to Appendix E.
B.4 Hierarchy examples
@@
Appendix C Full step count results
@@ Appendix D Algorithm
Appendix E Agent prompts

### v2

@@ Abstract
MLLM agents demonstrate potential for complex embodied tasks by retrieving multimodal task-relevant trajectory data. However, current retrieval methods primarily focus on surface-level similarities of textual or visual cues in trajectories, neglecting their effectiveness for the specific task at hand. To address this issue, we propose a novel method, MLLM As ReTriever (MART), which enhances the performance of embodied agents by utilizing interaction data to fine-tune an MLLM retriever based on preference learning, such that the retriever fully considers the effectiveness of trajectories and prioritize them for unseen tasks. We also introduce Trajectory Abstraction, a mechanism that leverages MLLMs’ summarization capabilities to represent trajectories with fewer tokens while preserving key information, enabling agents to better comprehend milestones in the trajectory. Experimental results across various environments demonstrate our method significantly improves task success rates in unseen scenes compared to baseline methods. This work presents a new paradigm for multimodal retrieval in embodied agents, by fine-tuning a general-purpose MLLM as the retriever to assess trajectory effectiveness. All the code for benchmark tasks, simulator modifications and the MLLM retriever is available at https://github.com/PKU-RL/MART.
@@ 4.1 Experimental Setup
Table 1: Environment complexity comparison. AI2-THOR contains more than four times the number of interactive objects per scene, and more complex object hierarchy, compared to LEGENT. E.g., AI2-THOR supports relationships such as “inside” (e.g., inside a microwave (Figure 10(a)), or inside an open container, like a sink (Figure 10(c))).
@@ 4.1.1 Environments
To validate the effectiveness of our method in various environments, we perform evaluations on multiple scenarios in two environments, AI2-THOR (Kolve et al., 2017) and LEGENT (Cheng et al., 2024). Unlike LLM agents, which use only text as input and simplify the action space by employing teleportation actions (e.g., “move directly to target”), we believe that MLLM agents should undertake more challenging tasks as they have access to visual input and are no longer limited to a ‘blind’ mode of operation. Therefore, both target environments are multimodal, whose observations are egocentric images, and both make use of fine-grained control actions.
@@
Table 4 and Figure 4 show a comparison between the two environments. Besides having many more objects in its scenes, AI2-THOR supports more complex object hierarchy relationships, such as “inside” (e.g., “inside a cabinet”, which require open/close interaction to complete a task, or “a cup inside a sink”, i.e., an open container), not supported in LEGENT  [^1] Although there are cabinets in LEGENT, they do not need to be opened/closed to complete tasks..
@@ 4.1.2 Task Settings
AI2-THOR. To better assess the fine-grained control ability of MLLM agents to complete real-world embodied tasks, we integrate characteristics of two AI2-THOR-based benchmarks – ALFWorld (Shridhar et al., 2021) and ALFRed (Shridhar et al., 2020) – and built an environment setting that is more suitable for the MLLM agent. Since tasks are long-horizon, we follow the method in ALFRed and apply task decomposition to divide them into sub-tasks before execution. Each sub-task is then provided to the agent, and it determines sub-task success based on environmental feedback by itself (details in Appendix F.2 and  F.3 ). Once a sub-task is successfully completed, the agent proceeds to the next sub-task. Task types include pick_and_place, pick_clean_then_place, pick_cool_then_place, and pick_heat_then_place. Completing these tasks requires dozens of steps of navigation, as well as interaction with objects. There are 45 tasks comprising a total of 260 sub-tasks in training set, and 28 tasks including 158 sub-tasks in testing set.
@@ 4.1.4 Task Evaluation
Average Steps represents the average number of steps the agent takes to complete a task. In AI2-THOR  it indicates the number of steps to complete a full task, and AS-Sub represents the step average to complete a sub-task. Notably, for failed cases, their steps are counted as the step limit.
@@ 4.1.5 Baselines
LLaVA-Plain (LP) is a pre-trained LLaVA with no modified head and no finetuning. We use the probability of special token generation to represent the score. Its input is the same as MART, and it is prompted to output only Yes/No tokens, and the final score is calculated based on the probability of token generation (more details and limitations in Appendix E.3).
@@ 4.5 Case Study
We present two case studies for more in-depth discussion of MART’s capabilities handling challenges of the MLLM agent setting. The first case (Figure 5) demonstrates the effective handling of a very long-horizon trajectory. Given a 73-step task trajectory – “navigate to DishSponge, put it in the Plate, and place the Plate on the Cabinet” – Trajectory Abstraction identifies 5 key milestones. MART achieves an 80% success rate with an average of 28 steps, while both the similarity-based method and the agent without memory reach only 40% success rate, averaging 69.6 and 74.6 steps, respectively. We provide more detailed and balanced case studies in Appendix D.
@@ 5 Conclusion
6 Discussions and Future Works
Our study has a few limitations. First, the MLLM’s restricted context window limits its ability to process multiple images simultaneously, restricting us to one-shot learning with single trajectory inputs. Future work will explore few-shot learning to combine skills from multiple trajectories for enhanced performance in complex tasks.
Second, the lack of a detailed ablation study will be addressed by evaluating each component’s contribution, including self-reflection mechanisms, prompt designs tailored to specific functions in action planning, and retrievers based on diverse base models.
Third, to ensure fair comparisons, we will construct large-scale, high-quality datasets specific to the household domain and fine-tune general feature extractors on these datasets to provide a more fair and direct comparison with similarity retrieval methods.
Fourth, model transfer stability across environments within the same domain needs validation. There are significant differences when comparing the household domain with the open-ended sandbox game domain and the web domain, including variations in task nature, operational frequency, object morphology, and textual input complexity. Therefore, we will train retrievers for the web, open-ended sandbox game, and household domains, respectively, to evaluate their generalization capability in novel environments and scenarios within each domain.
Lastly, the experimental scenarios in this study are somewhat limited, as we have only conducted experiments in the household domain. In future work, we plan to extend our approach to the web and open-ended sandbox game domains by collecting preference data through interactions and training retrievers for each domain. Subsequently, we will evaluate the generalization capabilities within the same domain, but in unseen environments and scenarios.
Acknowledgments
This work was supported by NSFC under Grant 62450001 and 62476008. The authors would like to thank the anonymous reviewers for their valuable comments and advice.
Appendix A Implementation details
In this section, we provide more implementation details about the model, training process and implementation pipeline.
A.1 Model and training details
We transform a generative language model (MLLM) into a trajectory scoring model by replacing the language model head with a Bradley-Terry score head. In particular, both the original language model head and our proposed Bradley-Terry scoring head are single-layer MLPs. However, there are notable differences: the language model head processes all hidden states as input and generates a probability distribution over the vocabulary for each token, facilitating token sequence generation through sampling. In contrast, our Bradley-Terry scoring head relies solely on the last non-zero hidden state as input and outputs a single floating-point score. Using this approach, our model generates only one new token at a time (i.e., by setting ‘max_new_tokens’ to 1). In comparison to conventional MLLM training, which generates hundreds or even thousands of new tokens per iteration, our model’s training is significantly more computationally efficient. All the code for the MLLM retriever model, training process, benchmark tasks and simulator modifications is available at https://github.com/PKU-RL/MART.
During training, we firstly fine-tune LLaVA to enable it to understand multiple images, details in E.1. After that, we replace the language model head with the Bradley-Terry score head, and fine-tune the model with lora. The parameter settings are listed in 6.
A.2 Details of implementation pipeline
In this section, we provide more details about the whole pipeline of implementation, including the data collection and retrieval in downstream tasks. The implementation pipeline is as follows:
  1. 1.
    Construct memory databases containing expert trajectories via the planner-based method, details in 3.2.
  2. 2.
    Collect the pairwise comparison data via interactive feedback to train the retriever model.
    1. (a)
      Specifically, for each task in the training set, we sample $K$ trajectories from the training memory $\mathcal{M^{\textrm{train}}}$, and feed them as prompts for MLLM agent to execute the embodied task respectively.
    2. (b)
      After that, based on the induced success rates of task execution, we can get the effectiveness of each trajectory for the embodied task. We can then obtain a partial order list based on success rate comparisons, producing $\binom{K}{2}$ pairs through pairwise comparison, where the trajectory with a higher success rate is treated as the positive item, and the one with a lower success rate as the negative item.
    3. (c)
      In this way, these preference pairs from interactive feedback are arranged as a positive-negative pair dataset $D$, which we use to fine-tune the MLLM according to the Bradley-Terry (Bradley & Terry, 1952) Reward Modeling loss to enhance its critiquing ability, as in Equation 1. Details are listed in Appendix 3.3 line 249-259.
  3. 3.
    Train the modified retriever model using the preference data. Details are listed in A.1.
  4. 4.
    Evaluation on unseen tasks. For each unseen task, we first retrieve trajectory with highest score for current task. Then the retrieved trajectory will be simplified through Trajectory Abstraction module. After that, the MLLM agent will execute the task with the help of abstraction of retrieved trajectory. Details are listed in Algorithm 1.
Appendix B Extension of related works
B.1 Embodied Grounding
Grounding is a critical challenge in embodied agents, referring to the alignment between the agent and its environment.
The grounding problem can be categorized into visual grounding and embodied grounding. Visual grounding (Lai et al., 2024; Kazemzadeh et al., 2014; Nagaraja et al., 2016) addresses the problem at the perception level by identifying the most relevant object or region in an image based on a language query, whereas embodied grounding focuses on the effects of actions on environmental dynamics and how an agent generates action sequences to accomplish a given task. Approaches to addressing embodied grounding can be categorized into the following types:
1. RL: Reinforcement learning (RL) trains an agent’s policy through interaction with the environment, making the agent inherently grounded in the environment, such as PPO (Schulman et al., 2017) and SAC (Haarnoja et al., 2018). However, RL typically requires extensive interaction with environments and often suffers from instability, making it unsuitable for MLLMs.
2. VLA: These methods focus on fine-tuning vision-language models (VLMs) using expert datasets collected from embodied environments, such as PaLM-E (Driess et al., 2023) and RT-2 (Brohan et al., 2023). These methods demand a significant amount of high-quality trajectory data for training.
3. LLM as Planner: These methods leverage Large Language Models (LLMs) or Multimodal Large Language Models (MLLMs) to generate high-level plans, which are then translated into executable action sequences by low-level controllers, such as LLM-Planner (Song et al., 2023) and P-RAG (Xu et al., 2024). A key limitation of these methods is their reliance on a predefined skill library, which restricts the scope of the agent’s capabilities. Besides, acquiring a skill library might require additional RL or IL training or prior knowledge about the environment (Lifshitz et al., 2023; Yuan et al., 2024).
4. Retrieval-Augmented MLLM Agent: This category involves integrating task trajectory data into the prompts provided to MLLMs, such as RAP (Kagaya et al., 2024). These trajectory data, rich in grounding information about the environment, enable agents to perform tasks effectively. Retrieval-augmented methods usually demonstrate greater sample efficiency compared to RL and VLA, thanks to the use of an explicit memory buffer. Our work falls into this category.
B.2 Multi-modal Information Retrieval
Appendix C Extension Experiments
We also evaluate our method on ReALFRED (Kim et al., 2025) environments, which provides realistic 3D-captured and multi-room scenes, as shown in Figure 7. The action space is similar with AI2-THOR, including fine-grained movement actions, e.g. ‘move ahead’, ‘turn left/right x degrees’, ‘look up/down’, and interactive atomic action including ‘pick up object A’, ‘put object A on/in object B’, ‘open/close object A’, ‘toggle on/off object A’. Since completing tasks often involves navigating across rooms, and the scenes within these rooms closely resemble real-world environments, this setting offers a more diverse and challenging scenario. The chosen task type is ‘pick_and_place’, which requires the agent to first navigate to the target object, pick it up, and then transport it to the designated location for placement. There are 30 tasks comprising a total 60 sub-tasks in training set, and 20 tasks including 40 sub-tasks in testing set.
Figure 7: Image examples of ReALFRED.
The experimental results demonstrate the effectiveness of our approach compared to the baselines. As shown in Table LABEL:tab:realfred_results, MART surpasses all baselines 10% in Success Rate, and reaches best performance across all metrics.
Table 5: Performance comparison of different methods in ReALFRED.
PA LP SL RAP MART
SR $\uparrow$ 0.25 0.20 0.26 0.27 0.37
SR-Sub $\uparrow$ 0.50 0.44 0.52 0.53 0.58
AS $\downarrow$ 101.70 87.93 89.79 87.89 78.48
AS-Sub $\downarrow$ 28.43 24.32 24.84 24.31 21.71
Appendix D Detailed Case Study
We present more detailed case studies, encompassing both success and failure cases for each method, along with simplified reasoning processes for clarity.
Figure 8: Comparison between unsuccessful and successful trajectory with MART.
Figure 9: Comparison between unsuccessful and successful trajectory with similarity-based retrieval.
In the first detailed case, as shown in Figure 8, the MART Retriever successfully retrieved a trajectory containing the target object’s location, while Trajectory Abstraction effectively compressed a 73-step trajectory into only 5 significant milestones, preserving crucial information. For the successful trial, the agent identified the target object’s location (the plate on the table) through the retrieved trajectory. After some exploration, it successfully found the target object and completed the task. For the unsuccessful trial, during exploration, the agent made mistakes (highlighted in purple) and failed to complete the task within the step limit.
In the second detailed case, as shown in Figure 9, the similarity-based retriever retrieved a trajectory that appeared similar but lacked useful information. As a result, the agent had to explore independently. By chance, the agent located the target object, but the overall success rate remained low.
Appendix E Experimental setup details
@@
In this appendix, we provide more low-level details on the implementation of MART experiments.
E.1 Multiple image input in LLaVA
@@
We utilized LLaVA’s single-image perception capabilities to enhance its multi-image perception. Specifically, we sampled images collected from the environments and employed the pre-trained LLaVA to describe the content of each image. We organized the responses into multi-image captioning data, which we then used to fine-tune LLaVA. After our verification, the fine-tuned LLaVA demonstrated the ability to perceive multiple input images. After fine-tuning with multiple image datasets, LLaVA serves as the basez model for our subsequent retriever training.
E.2 Parameter settings
@@
E.3 Details of LLava-Plain
Building on existing work (Asai et al., 2024) (Sun et al., 2023b) that employs LLM for text retrieval, we use the generation probability of a special token to represent the score for LLaVA-Plain. In detail, the effectiveness score $s_{i}$ is measured by the probability of LLaVA-Plain to generate the special token ‘Yes’ and ‘No’, as in Equation 2, where $p\left(\mathrm{Yes/No}\right)$ denoted the probability of LLaVA-Plain to generate Yes or No. I.
@@ E.3 Details of LLava-Plain
In detail, the prompts we use for LLaVA-Plain and MART are presented in prompt 1 of Appendix I.
Appendix F AI2-THOR environment specifics
F.1 Setting differences
@@
ALFWorld adapts ALFRED for LLM agents (i.e., text-only) by simplifying it. Firstly, it provides text feedback as observation, detailing objects in the agent’s field of view along with their corresponding IDs. Secondly, it simplifies the action space by replacing all navigation actions with the teleportation action ‘go to’ and composite high-level actions like ‘heat’, ‘clean’, and ‘cool’, each involving multiple atomic interactions. For example, the “cool object a” action is equivalent to: ‘open the refrigerator’, ‘put object A inside the refrigerator’, ‘close the refrigerator’, ‘open the refrigerator’, and ‘pick up the object A’. These modifications significantly reduce task difficulty, while allowing LLM agents to perform in ALFWorld.
@@
F.2 Task Decomposition
@@
F.3 Success Detection
In AI2-THOR, we provide the agent with the environment’s metadata and feedback in the Success Detection module. After each step is executed, the agent then determines whether the current sub-task is completed using a few-shot approach. For specific prompts, please refer to Appendix I.
F.4 Hierarchy examples
@@
Appendix G Full step count results
@@ Appendix H Algorithm
Appendix I Agent prompts